It takes 30 people 60 working days to build a small bridge. How many people are needed if the bridge is to be built in 40 working days, assuming that they all work at the same rate.
Answer by stanbon(75887)

It takes 30 people 60 working days to build a small bridge. How many people are needed if the bridge is to be built in 40 working days, assuming that they all work at the same
Days and # of people are inderectly related.
p = k/d
30 = k/60
k = 180
Equation
p= 180/40
p = 4.5 people



==========
Cheers,
Stan H.